OpenStudy (anonymous):

Solve for x: x^2 + 7x + 10 = 0 A.) -2, -5 B.) -2, 5 C.) 2, -5 D.) 2, 5

OpenStudy (anonymous):

a) : -2 , -5 :D

OpenStudy (anonymous):

thankya!

Parth (parthkohli):

@yashdhanuka I guess you have to explain your answer.

OpenStudy (anonymous):

uhh..yes but I think she didn't want the explanation because I just answered another quesiotn before :P

OpenStudy (anonymous):

i didn'tttt lol

Parth (parthkohli):

Wants are much more different than needs. Explain her the solution, but don't provide the answer.

OpenStudy (anonymous):

x^2 + 7x +10=0 x^2 + (5+2)x +10=0 x^2 + 5x+2x +10=0 x(x+5)+2(x+5)=0 x+2=0 or x+5=0 therefore x=-2,-5 :D :D